A window washer is standing on a scaffold supported by a vertical rope at each end. The scaffold weighs 200 N and is 2.90m long. What is the tension in each rope when the 700 N worker stands 1.90 m from one end?

smaller tension
larger tension
Answer #1

Tension in each rope due ro weight of scaffold=200/2=100 N

Tenaion in ropes due to worker's weight can be determined from moment equilibrium about the rope

From momemt equilibrium of scaffold about rope nearer to worker, we get

Tension in rope farther from worker=700*(2.9-1.9)/2.9=241.4 N

Tension in rope nearer to worker=700-241.4=458.6 N

Total tension in rope nearer the worker=458.6 +100=558.6 N

Total tension in rope farther from worker=100+241.4=341.4 N